Piper purchased one IBM May $100 call option at $6 and wrote one IBM May $110 call at $2.5.

a. What is the maximum potential profit of this strategy?

b. What is Piper’s profit if at option expiration the stock is trading at $105?

c. What is the maximum loss from the strategy?
Answer #1

rate positively ..

Ans 1 Gain on strategy
Profit on call buy = Max of ( price on expiry/sales - Strike price, 0) - Premium paid
Maximum gain on Call sale = Minimum of ( price on expiry/sales - Strike price, 0)+Premium on call written
Since strike price of call written is 110, therefore maximum gain will be at price of 110.
Profit on call buy = (110-100)-6
Profit on call sale = 0+2.5
Total gain = (110-100)-6+2.5
6.5
Ans 2 IF price is 105
Profit on call buy = (105-100)-6
Profit on call sale = 0+2.5
Total profit = (105-100)-6+2.5
1.5
Ans 3 Maximum loss will be when the call purchased is out of money means at or below $100
Loss on call buy = (100-100)-6
Profit on call sale = 0+2.5
Total profit = (100-100)-6+2.5
-3.5